# Heliometric stereo: a new frontier in surface profilometry

### v1

@@ VI The effects of masking, shadowing and multiple scattering
multiple scattering - any beam particle path that scatters from the sample more than once before detection.
@@
Masking, highlighted in area D of figure 7, causes dark regions in images, which will be reconstructed (equation (4)) as regions parallel to the detection direction. Shadowing results in areas of the sample where no information is gathered. Multiple scattering results in brighter regions of images, generally near vertical walls, recesses or similar features, such as in area C in figure 7, where a small recess at the edge of the embedded sphere results in enhanced contrast[24]. Bright regions caused by multiple scattering will also distort the solution to equation (4). the two spheres shown in figure 7 as our quantitative test of the impacts of masking and shadowing. The spheres provide a good test sample to quantify the effect of masking and shadowing on the method because both spheres have identical, and known, geometries, aside from sphere B having a larger aspect ratio because it is embedded in the substrate to a lesser degree. With masking and shadowing being geometrically defined as lines-of-sight between the sample and the detector or the incident aperture, respectively, when the aspect ratio of a feature increases the degree of masking and shadowing must also increase. This investigation is similar to that performed on simulated data in our previous work, where reconstruction accuracy was plotted as a function of the aspect ratio of features[26].
@@
For the lower aspect ratio sphere A, there is reasonably good accuracy across the sample, with only some small regions of poorer precision. In particular, the accuracy drops near the edge of the sphere, likely due to a crevice where the sphere is embedded into the surface, causing masking and multiple scattering. These errors are, however, locally confined, and do not cause errors away from the edges, similar to results found with simulated data[26]. Across the sphere itself the errors are asymmetric and match with the direction of the beam, but remain generally modest, generally reaching no more than double the average error.
@@
By contrast, while sphere B has the same basic shape as sphere A its increased aspect ratio results in more masking and shadowing. As can be seen in figure 9 the reconstructed surface lies noticeably higher towards the incident beam, and lower away from it, resulting in a visibly deformed reconstruction (figure 7). If we inspect quantitatively, we also find that the magnitude of the maximum deviation from the expected surface is also significantly greater than for sphere A, reaching and in percentage and shape, respectively. The greatest errors occur near the edge of the sphere, where there is non-negligible masking in all three micrographs, one of which is highlighted as feature D in figure 7, and where the surface of the sphere becomes approximately parallel to the beam.
Similar distortions were seen with simulated data when masking[26]; the simulated results suggest that by excluding masked regions from reconstruction more accurate shapes can be recovered, however, in order to exclude pixels and still perform a reconstruction more than 3 images are needed, which was not possible with the current experimental setup.
@@
Sphere B also demonstrates the effects of multiple scattering. On the flat substrate near the sphere there is a slight distortion in the reconstruction, the surface is higher than expected, e.g. in highlighted region E in figure 9. Here, the intensity of the helium micrographs has been increased in a similar manner to that shown in figure 8 (c), causing distortions in the reconstruction.
As, in most circumstances, multiple scattering only accounts for a small fraction of SHeM contrast these distortions have remained modest. However, multiple scattering is intrinsic to neutral atom microscopy and can only be controlled to a limited degree with instrument geometry[19], thus future applications of heliometric-stereo will need to consider it carefully.
@@
Shadowing, where the line-of-sight between incident beam and sample is blocked, is not present to a significant degree in the test samples considered here, however, it could lead to problems in reconstructions as parts of the sample could be missed. For a normal incidence configuration, such as that used in section IV, shadowing will occur at overhangs and vertical faces. For a non-normal incidence configuration, such as used in section V, vertical faces and overhangs can be captured if they face the beam, but regions facing away from the beam will be hidden. One approach to minimize the effect of shadowing would be to combining multiple reconstructions from different incidence conditions with the sample in different orientations, however, doing so would increase acquisition time.
Masking, which is intrinsic to neutral atom imaging, can be partially mitigated by changing the relative positions of the beam and the detectors: the larger the angular distance between the detector and beam, the greater the amount of masking[44]. An extreme optimization would be to place the detector in the same direction as the incident beam using an annular detector that surrounds the detector. The annular detection geometry is practicable experimentally with current polymer 3D printing methods used to create the pinhole plates for A-SHeM and B-SHeM. However, the annular detection geometry could only be used for one of the detectors so would either have to be combined with detectors that do exhibit masking.
@@
We include a reconstruction of a salt (sodium chloride) crystal in figure 10 using the same experimental setup as section V. Panels (a)-(c) show the helium micrographs used, panel (d) the resulting reconstruction and panel (e) an optical micrograph of the sample. The crystal has relatively low aspect ratio overall, but does posses a few vertical features at its edges that create strong shadowing or masking. Despite the local regions of shadowing and masking a reconstruction is generated that captures the overall topography of the sample well, highlighting the robustness of our method. By comparison to an optical microscope image (using a Keyence VHX-7000 microscope) we see that the heliometric stereo reconstruction reproduces the fine faceted structure on the surface of the crystal, including the cracks connecting the corners of the crystal, along with very fine ripples that appear parallel to the sides of the crystal forming concentric squares, which are of a size comparable to that of a single pixel in the SHeM micrographs. We note that these very fine features would not be recoverable with a single detector instrument as the lateral accuracy would be degraded by the image alignment procedure.
@@ VII Conclusion
Supplementary information
VIII Single detector sample SEM image
IX Imaging parameters - single detector
The acquisition time for micrographs in figure 4 was 3h 24min per micrograph. Pixel sizes of were used with a dwell time of . pixels were used for the micrograph.
The acquisition time for the micrograph in figure 3 was 7h 16min. Pixel sizes of were used with a dwell time of . pixels were acquired – the version presented is cropped.
X Imaging parameters - multiple detector
The acquisition time for all three micrographs needed for a single reconstruction, as shown in figure 7 for two spheres, was approximately 8 hours for images with dimensions pixels and dwell time per pixel.
The acquisition time for all three micrographs needed for the reconstruction of a salt crystal in figure 10 was approximately 10.5 hours for images with dimensions pixels and dwell time per pixel.
Per surface reconstruction, the effective speed-up using the multiple detector instrument is compared to the single detector implementation, due to both simultaneous image acquisition and the removal of the manual sample alignment procedure.
Acknowledgements
@@
The authors have no conflicts of interest to report.
@@
Supporting data packs containing both raw and processed data, and Supplementary Material can be found at http://doi.org/10.17863/CAM.107537 for results with the Cambridge A-SHeM and https://doi.org/10.17863/CAM.114465 for results with the Cambridge B-SHeM.

### v2

@@ VI The effects of masking, shadowing and multiple scattering
multiple scattering - any beam particle path that scatters more than once before detection.
It should be noted that the definitions of masking and shadowing used in the SHeM community are aligned with those used in optical, rather than electron, microscopy. An example is given by Linli Sun in their figure 2[46] for the case of optical microscopy. The convention was adopted upon the first formalization of these contrast mechanisms in SHeM by Fahy et al. [21].
@@
Masking, highlighted in figure 7 (f), causes dark regions in images, which will be reconstructed (equation (4)) as regions parallel to the detection direction. Shadowing results in areas of the sample where no information is gathered. Multiple scattering results in brighter regions of images, generally near vertical walls, recesses or similar features, such as in figure 7 (a), where a small recess at the edge of the embedded sphere results in enhanced contrast[26], and the elevated brightness on the substrate highlighted in figure 7 (e) as well as in smaller amounts on the sample itself. Bright regions caused by multiple scattering will also distort the solution to equation (4).
We use the two spheres shown in figure 7 as our quantitative test of the impacts of masking and shadowing. The spheres provide a good test sample to quantify the effect of masking and shadowing on the method because both spheres have identical, and known, geometries, aside from sphere B having a larger aspect ratio because it is embedded in the substrate to a lesser degree. With masking and shadowing being geometrically defined as lines-of-sight between the sample and the detector or the incident aperture, respectively, when the aspect ratio of a feature increases the degree of masking and shadowing must also increase. This investigation is similar to that performed on simulated data in our previous work, where reconstruction accuracy was plotted as a function of the aspect ratio of features[29].
@@
For the lower aspect ratio sphere A, there is reasonably good accuracy across the sample, with only some small regions of poorer precision. In particular, the accuracy drops near the edge of the sphere, likely due to a crevice where the sphere is embedded into the surface, causing masking and multiple scattering. These errors are, however, locally confined, and do not cause errors away from the edges, similar to results found with simulated data[29]. Across the sphere itself the errors are asymmetric and match with the direction of the beam, but remain generally modest, generally reaching no more than double the average error. The reconstruction also marginally underestimates the overall height of the sphere, recovering it at 94% of the expected height.
@@
By contrast, while sphere B has the same basic shape as sphere A its increased aspect ratio results in more masking and shadowing. As can be seen in figure 9 the reconstructed surface lies noticeably higher towards the incident beam, and lower away from it, resulting in a visibly deformed reconstruction (figure 7). If we inspect quantitatively, we also find that the magnitude of the maximum deviation from the expected surface is also significantly greater than for sphere A, reaching and in percentage and shape, respectively. The greatest errors occur near the edge of the sphere, where there is non-negligible masking in all three micrographs, one of which is highlighted in figure 7 (f), and where the surface of the sphere becomes approximately parallel to the beam. In the RMS errors in table 2 we also see that for sphere B, where masking is significant the error does not decrease when the reconstruction is allowed to scale (shape error), unlike for sphere A. A reduction in the error when scaling is allowed suggests that for sphere A the shape of the sample was well recovered, but that the height was underestimated. The lack of change (or negligible increase) in RMS error when the reconstruction is scaled suggest that there are significant errors in the shape of the sample, not merely a overall underestimation of the height. In the simulated results of Lambrick & Salvador Palau et al.[29] (figures 13/14) it was found that inclusion of (modest) multiple scattering induced an underestimation of the overall height, but that the shape was still recovered well; however, when the aspect ratio of samples was increased and masking appeared the shape was effected strongly. We believe the differences between the RMS and shape error in spheres A/B are due to the absence/presence of significant regions of masking, which sphere A does not posses and sphere B does. The simulated results suggest that by excluding masked regions from reconstruction more accurate shapes can be recovered, however, in order to exclude pixels and still perform a reconstruction more than 3 images are needed, which was not possible with the current experimental setup.
We find for both spheres that the overall height was recovered well, with 94% and 109% of the expected height recovered for spheres A and B respectively. This performance is slightly better than expected from simulations, which generally demonstrated a more significant underestimation of the overall height than we observe experimentally. Figure 10 gives the reconstructed surfaces and percentage error maps (same scale as figure 9). Here, as with the experimental data, we see a good quality reconstruction for sphere A, with the shape well preserved, but some distortion of the shape with sphere B. Unlike for the experimental data we underestimate the height of the sphere, recovering 85% and 76% of the expected height, as can be seen in the error maps. These values align with previous results where the dependence of the error on aspect ratio was explored[29]. The qualitatively different error maps between the simulated and the experimental data are interesting, and suggest that the ray-tracing simulations, which have a good track record of producing representative SHeM data[22, 12, 10, 26] may be incomplete for the current experimental setup.
The reconstruction process outlined by equations 2 allows the recovery of the ‘albedo factor’. While we do not attribute direct physical meaning to the recovered value we do observe rapid changes in the recovered value where deviant contrast features such as masking and strong multiple scattering arise. Figure 11 maps the recovered albedo factor for sphere’s A and B. In most areas the albedo varies slowly, however, strong changes occur near masking in sphere B, where we know that the relationship breaks down. For the test sample these maps of the albedo do not directly inform us of anything new, however, for more complicated samples where interpretation of the images is not as clear and the sample topography not known the albedo may become helpful in identifying regions of low accuracy reconstruction.
@@
Sphere B also demonstrates the effects of multiple scattering. On the flat substrate near the sphere there is a slight distortion in the reconstruction, the surface is higher than expected, e.g. in highlighted region in figure 9, which is likely to arise from brightening of the substrate, such as in figure 7 (e). It is possible the multiple scattering has also contributed to the reduced height of the final reconstruction. Here, the intensity of the helium micrographs has been increased in a similar manner to that shown in figure 8 (c), causing distortions in the reconstruction.
As, in most circumstances, multiple scattering only accounts for a small fraction of SHeM contrast these distortions have remained modest. There are specific sample geometries where the effects of multiple scattering are severe and unavoidable across much of the micrograph, such as high aspect ratio holes or troughs into a surface, however in most reported micrographs strong regions of multiple scattering tend to be localized to only part of the sample. An example of the obfuscation of the bottom of a narrow trough solely due to multiple scattering is shown by Lambrick et al.[26] in Figure 1. One can liken the outcome to the tip convolution problem in AFM and STM. Multiple scattering is intrinsic to neutral atom microscopy and can only be controlled to a limited degree with instrument geometry[21], thus future applications of heliometric-stereo will need to consider it carefully. The instrument geometry is especially relevant in pinhole microscopes designed to operate at short working distances, such as that presented by Witham et al.[4], where multiple scattering from the pinhole plate can become significant.
@@
Shadowing, where the line-of-sight between incident beam and sample is blocked, is not present to a significant degree in the test samples considered here, however, it could lead to problems in reconstructions as parts of the sample could be missed. For a normal incidence configuration, such as that used in section IV, shadowing will occur at overhangs and vertical faces. For a non-normal incidence configuration, such as used in section V, vertical faces and overhangs can be captured if they face the beam, but regions facing away from the beam will be hidden. One approach to minimize the effect of shadowing would be to combine multiple reconstructions from different incidence conditions with the sample in different orientations, however, doing so would increase acquisition time.
Masking, which is intrinsic to neutral atom imaging, can be partially mitigated by changing the relative positions of the beam and the detectors: the larger the angular distance between the detector and beam, the greater the amount of masking[47]. Both implementations in the current work already attempt to keep the angular distance between detector and beam modest, although it was found that the desire to reduce the angular distance conflicts with the limited space around the sample in both single and multiple detector SHeM, and in practice some compromise had, and will have, to be made in the design. For future instruments or modifications intended for the use of heliometric stereo, attempts to reduce masking must be kept in consideration alongside other requirements of the instrument.
@@
We include a reconstruction of a salt (sodium chloride) crystal in figure 12 using the same experimental setup as section V. Panels (a)-(c) show the helium micrographs used, panel (d) the resulting reconstruction and panel (e) an optical micrograph of the sample. The crystal has relatively low aspect ratio overall, but does posses a few vertical features at its edges that create strong shadowing or masking. Despite the local regions of shadowing and masking a reconstruction is generated that captures the overall topography of the sample well, highlighting the robustness of our method. By comparison to an optical microscope image (Keyence VHX-7000 microscope) we see that the heliometric stereo reconstruction reproduces the fine faceted structure on the surface of the crystal, including the cracks connecting the corners of the crystal, along with very fine ripples that appear parallel to the sides of the crystal forming concentric squares, which are of a size comparable to that of a single pixel in the SHeM micrographs. We note that these very fine features would not be recoverable with a single detector instrument as the lateral accuracy would be degraded by the image alignment procedure.
We have demonstrated that heliometric stereo can be applied successfully both with a single detector SHeM, section IV, and with a multiple detector SHeM, section V. The primary advantage explored thus far for the multiple detector approach is the speed up in measurement time, with a speed up of directly achieved in the presented results, and speed ups of roughly achievable with the presented setups depending on specific parameters used. An apparent drawback of the multiple detector approach presented here, where a non-normal incidence angle is used, is the limited number of micrographs available, 3 instead of 5. In general more micrographs increase the SNR, but more importantly, they can be used to get around masking issues[29] by allowing certain data regions to be excluded from the reconstruction while still having sufficient data to solve equations 2. In the multiple detector case any incidence angle can be used, while the single detector case is limited to normal incidence. As the reconstruction method performs poorly for samples areas parallel to the beam, varying the incidence angle for the specific sample of interest could be a benefit for some systems. Alternatively if a normal incidence beam were to be used for a multiple detector system a hybrid approach can be employed, where rotations are used to multiply the number of detector, e.g. a single rotation could provide 6 effective detectors with a 3 detector microscope setup. The hybrid approach would yield more modest speed ups, but would provide more data points and hence likely a high reconstruction quality than either approach presented in the current work.
As well as developments of the experimental hardware to overcome the limitations imposed by masking and multiple scattering, we can also envisage future improvements to the computational methods. For example, stereo-photogrammetry or masks in micrographs can be used to give accurate dimensions for specific features[25, 26], these could be used to calibrate and improve the accuracy of heliometric-stereo reconstructions where height of features can be underestimated (such as the larger sphere in figure 7). In addition, more involved iterative methods could be used to reduce the uncertainties arising from multiple scattering, an issue that may become even more relevant if microscopes move to shorter working distances to improve spatial resolution. By using tools developed to predict contrast in SHeM that account for multiple scattering[22] the multiple scattering intensities from the initially reconstructed surface could be predicted, those predictions could then be used to inform/correct a second (or further) reconstruction(s). Potentially more advanced parametric, or even non-parametric[48], models for the BRDF could be used, loosening the hard assumption of diffuse cosine-like scattering, however, adding more parameters would entail the collection of a greater number of micrographs per sample.
@@ VII Conclusion
Supplementary Material
Supplementary Material
SI Single detector sample SEM image
SII Imaging parameters - single detector
The total acquisition time for the micrographs in figure 4 was 14 hours and 53 minutes (2x images with acquisition time of 3 hours 25 minutes and 3x images with acquisition time of 2 hours 41 minutes). Pixel sizes of were used with a dwell time of for 2x images and for 3x images. pixels were used for the micrograph.
SIII Imaging parameters - multiple detector
The acquisition time for all three micrographs needed for a single reconstruction of a sphere, as shown in figure 7 (same parameters used for both spheres), was 5 hours and 14 minutes for images with dimensions pixels and dwell time per pixel.
The acquisition time for all three micrographs needed for the reconstruction of a salt crystal in figure 10 was 8 hours and 8 minutes for images with dimensions pixels and dwell time per pixel.
SIV Error calculation areas
For the calculation of the RMS and shape error most of the flat substrate was not included in calculations. A circular region centered on the sphere was used, these circles are drawn over the percentage error in figure S2.
Acknowledgments
@@
The authors declare the following financial interests/personal relationships which may be considered as potential competing interests: Sam Lambrick reports a relationship with lonoptika Ltd. that includes: consulting or advisory. David Ward reports a relationship with lonoptika Ltd. that includes: consulting or advisory. The other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@
Supporting data packs containing both raw and processed data are available. Data for results with the Cambridge A-SHeM (figures 4 and 5) can be found at http://doi.org/10.17863/CAM.107537 . Data for figure 12 (and some of the data for figure 7) can be found athttps://doi.org/10.17863/CAM.114465. A data pack for figures 7, 9, 10, 11, and the optical data in figure 12 can be found at https://doi.org/10.5281/zenodo.15548542.